# Cycling Race Time Prediction: A Personalized Machine Learning Approach Using Route Topology and Training Load

### v1

@@ Abstract
Predicting cycling duration for a given route is essential for training planning and event preparation. Existing solutions rely on physics-based models that require extensive parameterization, including aerodynamic drag coefficients and real-time wind forecasts, parameters impractical for most amateur cyclists. This work presents a machine learning approach that predicts ride duration using route topology features combined with the athlete’s current fitness state derived from training load metrics. The model learns athlete-specific performance patterns from historical data, substituting complex physical measurements with historical performance proxies. We evaluate the approach using a single-athlete dataset (N=96 rides) in an N-of-1 study design. After rigorous feature engineering to eliminate data leakage, we find that Lasso regression with Topology + Fitness features achieves MAE=6.60 minutes and R²=0.922. Notably, integrating fitness metrics (CTL, ATL) reduces error by 14% compared to topology alone (MAE=7.66 min), demonstrating that physiological state meaningfully constrains performance even in self-paced efforts. Progressive checkpoint predictions enable dynamic race planning as route difficulty becomes apparent.
@@ 2.1.  Cycling Performance Prediction
The prediction of cycling performance has traditionally relied on physics-based models. Di Prampero et al. [diprampero1979equation] established the foundational equation of motion for cyclists, relating power output to resisting forces: aerodynamic drag, rolling resistance, gravitational force on gradients, and drivetrain losses. Martin et al. [martin1998validation] validated this mathematical model for road cycling power, demonstrating accurate prediction of power requirements under controlled conditions. This equation forms the basis of commercial tools that predict race times and optimal pacing strategies.
@@
Machine learning has been applied extensively to sports prediction problems [chen2020machine]. In cycling, researchers have used gradient boosting methods to predict race outcomes, leveraging historical results, course profiles, and athlete rankings. Deep learning approaches, including LSTM networks, have modeled sequential performance data to capture temporal patterns in athlete form. Jobson et al. [jobson2009prediction] demonstrated the utility of cycling training data analysis for performance modeling, while Menaspà et al. [menaspa2017cycling] characterized the physical demands of professional cycling using power meter data.
@@ 2.3.  Training Load and Fitness Modeling
Training Stress Score (TSS) quantifies the physiological load of individual workouts based on intensity and duration relative to threshold power [coggan2019tss]. Chronic Training Load (CTL) represents the exponentially weighted average of TSS over approximately 42 days, serving as a proxy for fitness. Acute Training Load (ATL) uses a shorter time constant (approximately 7 days) to capture recent fatigue. Training Stress Balance (TSB), calculated as CTL minus ATL, indicates an athlete’s freshness and theoretical readiness to perform.
@@ 3.1.  Data Sources
The data for this study comes from a single amateur cyclist’s training records, collected over multiple years through standard cycling equipment. All data is sourced from Intervals.icu, a training analysis platform that aggregates data from cycling computers and power meters.
@@ 3.1.1. Data Platform
Raw data is processed through a medallion data architecture. The ingestion layer (bronze) retrieves data from the Intervals.icu API; a cleaning pipeline (silver) applies type coercion, schema validation, and unit standardization. The platform runs daily, providing clean, consistently formatted data for analysis.
@@ 3.1.3. Wellness and Training Load
Daily wellness records capture physiological state: body weight, Chronic Training Load (CTL), Acute Training Load (ATL), and Training Stress Balance (TSB). These are computed using established formulas based on historical power data and reflect the athlete’s state at the start of each day.
@@ 3.2.  Feature Engineering
rolling_*d_hr_z*_hours $H_{w,z}$ (Eq. 6) 5 HR zones $\times$ 4 windows
@@ 3.2.1. Route Topology Features
where $d$ is segment length in meters and $g$ is average gradient in percentage points (e.g., 8 for 8%), yielding an HC score of 80,000 for a 10km climb at 8%. Scores above 80,000 are Hors Catégorie (HC); 64k, 32k, 16k, 8k for Cat 1–4; >>1,500 for uncategorized climbs. We additionally compute Tour de France scores ($g^{2}\times d$) as a separate feature. From detected climbs we derive: count by category (HC, Cat 1–4), total and maximum scores, average gradient weighted by distance, climb density (climbs per km), and length metrics (average, maximum, total climb distance). Short climbs (1–5 min) tax anaerobic capacity; sustained climbs (>>10 min) depend on aerobic power [leo2022power].
Gradient Variability.
@@
Variable terrain causes 26% greater fatigue than constant grades at equal average power [leo2022durability]. This occurs because gradient variability forces repeated anaerobic efforts during accelerations on steep sections, prevents steady-state metabolic adaptation, and requires constant power adjustments that increase neuromuscular fatigue. We compute a punchiness score:
@@
Recovery and Technicality.
After significant climbs, flat sections allow W (anaerobic work capacity) reconstitution with time constants of 377–580s [skiba2015intramuscular]. We compute recovery distance as flat terrain within 500m of climb ends:
@@
These sections require sustained attention despite low power output, limiting recovery [abbiss2008describing].
@@
Elevation Profile Shape.
@@
Sharp Turns and Turn Density.
Sharp turns require braking, gear changes, and power surges to accelerate out, disrupting rhythm and accumulating fatigue [padilla2000scientific]. We detect sharp turns where bearing change exceeds $45^{\circ}$:
@@
Additional Route Metrics.
@@ 3.2.2. Athlete State Features
Training intensity distribution predicts performance [sanders2017methods]. Following Banister’s impulse-response model [banister1991modeling] and Seiler’s time-in-zone methodology [seiler2006quantifying], we compute rolling zone hours:
@@
HR zones: Z0–Z4 (5 zones) $\times$ 4 windows = 20 features
@@
    Topology + Fitness + Zones (79 features): Adds rolling zone hours (7d, 14d, 30d, 60d windows across power and HR zones). Captures recent training load distribution.
@@
  • Best accuracy: Use Topology + Fitness model (31 features) when historical training data is available. Requires athlete’s current CTL, ATL, TSB, and ramp rate from prior weeks of training.
@@ 3.4.1. Model Candidates
We evaluate three model families: (1) Regularized linear models (Ridge, Lasso, ElasticNet) with L1/L2 penalties to prevent overfitting, (2) Random Forest with Bayesian hyperparameter optimization via Optuna [akiba2019optuna], and (3) Baseline models (mean, median, simple linear regression on distance + elevation) to establish prediction difficulty.
@@ 3.4.2. Nested Cross-Validation
Outer Loop (Evaluation):
@@
Inner Loop (Hyperparameter Tuning):
@@ 4.1.  Dataset
The dataset comprises 96 outdoor cycling activities recorded between March 2023 and October 2025 from a single amateur cyclist (male, FTP 232W). Activities were classified as free rides (N=92) or races (N=4) using duration deviation from expected time. Table 2 summarizes route and performance characteristics.
@@ 4.3.  Error Analysis
To understand which route characteristics drive predictions, we analyze feature importance using SHAP (SHapley Additive exPlanations) [lundberg2017unified] values on the best model (Lasso with topology features).
@@ 6.1.1. Topology Sets the Baseline, Fitness Refines the Prediction
Route topology is the primary driver of duration, explaining over 88% of the variance in our dataset. Features like punchiness (gradient variability) capture difficulty beyond simple elevation gain [leo2022durability]. However, the integration of fitness metrics (CTL, ATL) provides a critical 14% improvement in accuracy (MAE reduced from 7.66 to 6.60 min).
@@ 6.1.2. Fitness Improves Prediction, but Interpretation Requires Caution
Despite this limitation, the aggregate contribution of fitness features is valid: including CTL, ATL, and related metrics consistently reduces prediction error across cross-validation folds. Future work with larger datasets or experimental designs (e.g., repeated measurements under controlled fitness variations) could enable more reliable causal inference for training periodization guidance.
@@ 6.1.5. N-of-1 Study Design Justification
Single-athlete studies are appropriate when: (1) High inter-individual variability: Cycling performance depends on FTP, VO2max, lactate threshold, W’, body composition, and skill—parameters varying 2-5x across amateur cyclists. Population models would require extensive physiological testing (lab VO2max, power profiling) impractical for the target use case. (2) Personalized prediction goal: This work targets individual athletes seeking “How long will this route take ME?” rather than “How long for an average cyclist?” Personalized models learn athlete-specific strengths (climbing vs. flat speed) and pacing tendencies. (3) Data availability: Modern cyclists generate thousands of GPS activities, providing abundant within-subject data without requiring sparse multi-athlete cohorts.
Similar N-of-1 designs have proven effective in chronobiology, nutrition, and clinical medicine [kravitz2014design]. The key is sufficient within-subject variation in predictor variables (route characteristics) and outcome (duration), which our dataset provides.
@@ 6.2.  Limitations
Environmental Factors Excluded.
Weather conditions (wind, temperature, precipitation), road surface quality, and traffic are not captured. Wind alone can alter duration by 10-30% on exposed routes. GPS data lacks environmental sensors; integrating weather API data is future work.
Pacing Assumptions.
@@
GPS Accuracy.
Elevation data from barometric altimeters has $\pm$5m precision, while GPS elevation can drift $\pm$20m. Systematic errors in climb detection or gradient calculation could propagate to predictions, though high-quality devices mitigate this limitation.
@@ 6.3.  Future Work
Multi-Athlete Validation.
@@
Real-Time Pacing Guidance.
@@
Segment-Level Predictions.

### v2

@@ Abstract
Predicting cycling duration for a given route is essential for training planning and event preparation. Existing solutions rely on physics-based models that require extensive parameterization, including aerodynamic drag coefficients and real-time wind forecasts, parameters impractical for most amateur cyclists. This work presents a machine learning approach that predicts ride duration using route topology features combined with the athlete’s current fitness state derived from training load metrics. The model learns athlete-specific performance patterns from historical data, substituting complex physical measurements with historical performance proxies. We evaluate the approach using a single-athlete dataset (N=96 rides) in an N-of-1 study design. After rigorous feature engineering to eliminate data leakage, we find that Lasso regression with Topology + Fitness features achieves MAE=6.60 minutes and R²=0.922. Notably, integrating fitness metrics (Chronic Training Load [CTL], Acute Training Load [ATL]) reduces error by 14% compared to topology alone (MAE=7.66 min), demonstrating that physiological state meaningfully constrains performance even in self-paced efforts. Progressive checkpoint predictions enable dynamic race planning as route difficulty becomes apparent.
@@ 2.1.  Cycling Performance Prediction
The prediction of cycling performance has traditionally relied on physics-based models. Di Prampero et al. [6] established the foundational equation of motion for cyclists, relating power output to resisting forces: aerodynamic drag, rolling resistance, gravitational force on gradients, and drivetrain losses. Martin et al. [12] validated this mathematical model for road cycling power, demonstrating accurate prediction of power requirements under controlled conditions. This equation forms the basis of commercial tools that predict race times and optimal pacing strategies.
@@
Machine learning has been applied extensively to sports prediction problems [5]. In cycling, researchers have used gradient boosting methods to predict race outcomes, leveraging historical results, course profiles, and athlete rankings. Deep learning approaches, including LSTM networks, have modeled sequential performance data to capture temporal patterns in athlete form. Jobson et al. [7] demonstrated the utility of cycling training data analysis for performance modeling, while Menaspà et al. [13] characterized the physical demands of professional cycling using power meter data.
@@ 2.3.  Training Load and Fitness Modeling
Training Stress Score quantifies the physiological load of individual workouts based on intensity and duration relative to threshold power [3]. Chronic Training Load (CTL) represents the exponentially weighted average of Training Stress Score over approximately 42 days, serving as a proxy for fitness. Acute Training Load (ATL) uses a shorter time constant (approximately 7 days) to capture recent fatigue. Training Stress Balance (TSB), calculated as CTL minus ATL, indicates an athlete’s freshness and theoretical readiness to perform.
@@ 3.1.  Data Sources
The data for this study come from a single amateur cyclist’s training records, collected over multiple years through standard cycling equipment. All data are sourced from Intervals.icu, a training analysis platform that aggregates data from cycling computers and power meters.
@@ 3.1.1. Data Platform
Raw data are processed through a medallion data architecture. The ingestion layer (bronze) retrieves data from the Intervals.icu API; a cleaning pipeline (silver) applies type coercion, schema validation, and unit standardization. The platform runs daily, providing clean, consistently formatted data for analysis.
@@ 3.1.3. Wellness and Training Load
Daily wellness records capture physiological state: body weight, CTL, ATL, and TSB. These are computed using established formulas based on historical power data and reflect the athlete’s state at the start of each day.
@@ 3.2.  Feature Engineering
rolling_*d_hr_z*_hours $H_{w,z}$ (Eq. 6) 5 heart rate zones $\times$ 4 windows
@@ 3.2.1. Route Topology Features
where $d$ is segment length in meters and $g$ is average gradient in percentage points (e.g., 8 for 8%), yielding an HC score of 80,000 for a 10km climb at 8%. Scores above 80,000 are Hors Catégorie (HC); 64k, 32k, 16k, 8k for Cat 1–4; >>1,500 for uncategorized climbs. We additionally compute Tour de France scores ($g^{2}\times d$) as a separate feature. From detected climbs we derive: count by category (HC, Cat 1–4), total and maximum scores, average gradient weighted by distance, climb density (climbs per km), and length metrics (average, maximum, total climb distance). Short climbs (1–5 min) tax anaerobic capacity; sustained climbs (>>10 min) depend on aerobic power [10].
Gradient Variability
@@
Variable terrain causes 26% greater fatigue than constant grades at equal average power [9]. This occurs because gradient variability forces repeated anaerobic efforts during accelerations on steep sections, prevents steady-state metabolic adaptation, and requires constant power adjustments that increase neuromuscular fatigue. We compute a punchiness score:
@@
Recovery and Technicality
After significant climbs, flat sections allow W (anaerobic work capacity) reconstitution with time constants of 377–580s [17]. We compute recovery distance as flat terrain within 500m of climb ends:
@@
These sections require sustained attention despite low power output, limiting recovery [1].
@@
Elevation Profile Shape
@@
Sharp Turns and Turn Density
Sharp turns require braking, gear changes, and power surges to accelerate out, disrupting rhythm and accumulating fatigue [14]. We detect sharp turns where bearing change exceeds $45^{\circ}$:
@@
Additional Route Metrics
@@ 3.2.2. Athlete State Features
Training intensity distribution predicts performance [15]. Following Banister’s impulse-response model [4] and Seiler’s time-in-zone methodology [16], we compute rolling zone hours:
@@
Heart rate zones: Z0–Z4 (5 zones) $\times$ 4 windows = 20 features
@@
    Topology + Fitness + Zones (79 features): Adds rolling zone hours (7d, 14d, 30d, 60d windows across power and heart rate zones). Captures recent training load distribution.
@@
  • Best accuracy: Use Topology + Fitness model (31 features) when historical training data are available. Requires athlete’s current CTL, ATL, TSB, and ramp rate from prior weeks of training.
@@ 3.4.1. Model Candidates
We evaluate three model families: (1) Regularized linear models (Ridge, Lasso, ElasticNet) with L1/L2 penalties to prevent overfitting, (2) Random Forest with Bayesian hyperparameter optimization via Optuna [2], and (3) Baseline models (mean, median, simple linear regression on distance + elevation) to establish prediction difficulty.
@@ 3.4.2. Nested Cross-Validation
Outer Loop (Evaluation)
@@
Inner Loop (Hyperparameter Tuning)
@@ 4.1.  Dataset
The dataset comprises 96 outdoor cycling activities recorded between March 2023 and October 2025 from a single amateur cyclist (male, Functional Threshold Power 232W). Activities were classified as free rides (N=92) or races (N=4) using duration deviation from expected time. Table 2 summarizes route and performance characteristics.
@@ 4.3.  Error Analysis
To understand which route characteristics drive predictions, we analyze feature importance using SHAP (SHapley Additive exPlanations) [11] values on the best model (Lasso with topology features).
@@ 6.1.1. Topology Sets the Baseline, Fitness Refines the Prediction
Route topology is the primary driver of duration, explaining over 88% of the variance in our dataset. Features like punchiness (gradient variability) capture difficulty beyond simple elevation gain [9]. However, the integration of fitness metrics (CTL, ATL) provides a critical 14% improvement in accuracy (MAE reduced from 7.66 to 6.60 min).
@@ 6.1.2. Fitness Improves Prediction, but Interpretation Requires Caution
Despite this limitation, the aggregate contribution of fitness features is valid: including CTL, ATL, and related metrics consistently reduce prediction error across cross-validation folds. Future work with larger datasets or experimental designs (e.g., repeated measurements under controlled fitness variations) could enable more reliable causal inference for training periodization guidance.
@@ 6.1.5. N-of-1 Study Design Justification
Single-athlete studies are appropriate when: (1) High inter-individual variability: Cycling performance depends on Functional Threshold Power, VO2max, lactate threshold, W’, body composition, and skill—parameters varying 2-5x across amateur cyclists. Population models would require extensive physiological testing (lab VO2max, power profiling) impractical for the target use case. (2) Personalized prediction goal: This work targets individual athletes seeking “How long will this route take ME?” rather than “How long for an average cyclist?” Personalized models learn athlete-specific strengths (climbing vs. flat speed) and pacing tendencies. (3) Data availability: Modern cyclists generate thousands of GPS activities, providing abundant within-subject data without requiring sparse multi-athlete cohorts.
Similar N-of-1 designs have proven effective in chronobiology, nutrition, and clinical medicine [8]. The key is sufficient within-subject variation in predictor variables (route characteristics) and outcome (duration), which our dataset provides.
@@ 6.2.  Limitations
Environmental Factors Excluded
Weather conditions (wind, temperature, precipitation), road surface quality, and traffic are not captured. Wind alone can alter duration by 10-30% on exposed routes. GPS data lack environmental sensors; integrating weather API data are future work.
Pacing Assumptions
@@
GPS Accuracy
Elevation data from barometric altimeters have $\pm$5m precision, while GPS elevation can drift $\pm$20m. Systematic errors in climb detection or gradient calculation could propagate to predictions, though high-quality devices mitigate this limitation.
@@ 6.3.  Future Work
Multi-Athlete Validation
@@
Real-Time Pacing Guidance
@@
Segment-Level Predictions